[-4]^3 and [-4^3] ,can you explain how having exponent outside the parentheses and inside parenteses works. example please. many thanks andy
Answer by sudhanshu_kmr(1152)


(-4)^3 = -64 (-4^3) = -64
value will be different if exponent value is an even number.....
(-4)^2 = 16 but (-4^2) = -16

if exponent value on a negative value is an even number, it will become
positive...
for example...
(-2)^4 = 16 , (-2)^5 = 32
(-2)^3 = (-2^3) = -8